OpenStudy (anonymous):

Find the slope and the y-intercept of the graph of the equation y-9x=0

jimthompson5910 (jim_thompson5910):

y-9x=0 y-9x+9x=0+9x y = 9x y = 9x + 0 I'll let you finish up

OpenStudy (anonymous):

Rewrite as y = mx + b where m is the slope and b is the y-intercept. So add 9x to each side of the equation: y = 9x + 0 and solve for m and b.